OpenStudy (anonymous):

A line has a slope of 2/3 and contains the points (x, 9) and (3, 7). The value of x is A.0 B.3 C.6

OpenStudy (anonymous):

\[(y _{2} - y _{1} )/(x _{2} - x _{1} )=m \] \[(7-9 )/(3- x) =2/3 \] \[-6=6-2x\] \[2x=12\] \[x=6 \]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

No problem